# ../../Bibliographies/main.bib

### v1

@@ 3.2. Well-Foundedness
Since $\Omega^{(0)}>\Omega^{(1)}>\cdots$, the ordinal terms are certainly not well-founded on their face. This is because our notation looks at ordinal terms “from above”, fixing $\Omega^{(0)}$ to be the highest cardinality. We should instead compare “from below”, demanding the ordinals share the same ground.
@@
We take the view that shifting the indices does not really change the meaning of an ordinal term.
@@
We define $[\alpha]_{g}=\{\alpha^{\geq 0}_{+m}\mid m\in\mathbb{Z},\overline{\mathbb{FC}}(\alpha)\leq-m\}$.
A minor complication is that we typically can’t raise the ground—in the ordinal term $\Omega^{(0)}\#\Omega^{(1)}\#\vartheta(\Omega^{(1)}\#\Omega^{(2)})$, trying to lower the indices by $1$ (so the ground would be 00) would create problems inside the $\vartheta$. Instead, we have to compare ordinals by lowering the ground until they match.
Definition 3.8.
We say $[\alpha]_{g}<[\beta]_{g}$ if for some (equivalently, any) $m,n\in\mathbb{N}$ large enough that either $\mathbb{FC}(\beta)=-\infty$ or $\mathbb{FC}(\alpha)=-\infty$ or $\mathbb{G}(\alpha^{\geq 0}_{+m})=\mathbb{G}(\beta^{\geq 0}_{+n})$, we have $\alpha^{\geq 0}_{+m}<\beta^{\geq 0}_{+n}$.
Each equivalence class $[\alpha]_{g}$ has a distinguished element $\alpha^{*}=\left\{\begin{array}[]{ll}\alpha&\text{if }\mathbb{FC}(\alpha)=-\infty\\ \alpha^{\geq 0}_{+\mathbb{FC}(\alpha)}&\text{otherwise}\end{array}\right..$ That is, we arrange to have the largest term appearing be $\Omega^{(0)}$.
@@
The ordinal terms of formal cardinality $\infty$- ∞ are well-founded.
@@
We will define two collections of sets $\mathrm{Acc}_{n}\subseteq M_{n}$ for $n\in\{-\infty\}\cup\mathbb{N}$. We let $M_{-\infty}$ be all ordinal terms of formal cardinality $\infty$- ∞.
Let $n\in\mathbb{N}$. If we have defined $\mathrm{Acc}_{i},M_{i}$ for $i<n$ and $S\subseteq\bigcup_{i<n}M_{i}$ then we write $S\subseteq\mathrm{Acc}$ if, for each $i<n$, $S\cap M_{i}\subseteq\mathrm{Acc}_{i}$.
We set $M_{n}$ to be the set of ordinal terms $\alpha$ with $\mathbb{FC}(\alpha)=0$ and $\mathbb{G}(\alpha)=n$ such that $\{\beta^{*}\mid\beta\in K^{\geq 0}\alpha\}\subseteq\mathrm{Acc}$. Note that if $\beta\in K^{\geq 0}$ then we have $\mathbb{G}(\beta)\geq\mathbb{G}(\alpha)$ while $\mathbb{FC}(\beta)<\mathbb{FC}(\alpha)$, so, since $\mathbb{FC}(\beta^{\geq 0}_{+\mathbb{FC}(\beta)})\in\{0,-\infty\}$, we have $\mathbb{G}(\alpha)<n$. We let $\mathrm{Acc}_{n}$ be the well-founded part of $M_{n}$.
[…]
We claim that for $n>-\infty$ and $\alpha\in\mathrm{Acc}_{n}$, $(\vartheta\alpha)^{*}\in\mathrm{Acc}_{\mathbb{G}((\vartheta\alpha)^{*})}$. We prove this by induction on $\mathrm{Acc}_{n}$, so consider some $\alpha\in\mathrm{Acc}_{n}$ so that, for all $\beta<\alpha$ in $\mathrm{Acc}_{n}$, $(\vartheta\beta)^{*}\in\mathrm{Acc}_{\mathbb{G}((\vartheta\beta)^{*})}$. Let $m=\mathbb{G}((\vartheta\alpha)^{*})$.
If $\gamma$ is $\Omega^{(0)}$, this is immediate since $\Omega^{(0)}$ is the smallest element of formal cardinality 00. If $\gamma=\#\{\gamma_{i}\}$ or $\gamma=\omega^{\gamma^{\prime}}$ then this follows by standard arguments and the side inductive hypothesis.
So suppose $\gamma=\vartheta\gamma^{\prime}$ note that $\gamma=\gamma^{*}$ since $\gamma\in\mathrm{Acc}_{m}$. We have two possibilities. If $\gamma^{\prime}<\alpha$ then, by the main inductive hypothesis, $(\vartheta\gamma^{\prime})^{*}\in\mathrm{Acc}_{m}$. Otherwise there is some $\delta\in K^{\geq 0}\alpha$ with $\gamma\leq\delta$, and by assumption, $\delta^{*}\in\mathrm{Acc}_{m}$, so $\gamma\in\mathrm{Acc}_{m}$ as well.
Finally, we show by induction on terms that every term $\alpha$ with formal cardinality in $\{0,-\infty\}$ belongs to $\mathrm{Acc}_{\mathbb{G}(\alpha)}$. Again the $+$ and $\omega^{\cdot}$ cases are standard, and the case $\alpha=\vartheta\alpha^{\prime}$ follows immediately from the previous step and the induction hypothesis. ∎
@@ 3.3. Variables
Finally, we should extend the system with a variable so that we can check the Key Lemma. We extend the notation system by variables $v^{(J)}$, where we interpret $v^{(J)}$ as smaller than $\Omega^{(J)}$ but larger than anything of formal cardinality $-(J+1)$.
@@
  • $(\vartheta\alpha)[v\mapsto^{J}\beta]=\vartheta(\alpha[v\mapsto^{J+1}\beta])$,
  • $w^{(J^{\prime})}[v\mapsto^{J}\beta]=\left\{\begin{array}[]{ll}\beta^{\geq 0}_{+J}&\text{if }v=w\\ w^{(J^{\prime})}&\text{if }v\neq w\end{array}\right.$.
We should only consider substituting into variables which appear precisely at the top level—that is, we want to substitute into things like $v^{(0)}\#\vartheta v^{(1)}$.
@@
  • $\alpha=\vartheta\alpha^{\prime}$ and $v$ is $J+1$-substitutable in $\alpha^{\prime}$,
@@
    If $\alpha<\beta$ and $K^{\geq 0}\alpha<\vartheta\beta$ then $\vartheta\alpha<\vartheta\beta$.
@@
    If $\alpha<\beta$, $\gamma<\beta$, $v$ is 00-substitutable in $\gamma$, $K^{\geq 0}\alpha<\vartheta\beta$, and $K^{\geq 0}\gamma<\vartheta\beta$, then $\vartheta(\gamma[v\mapsto\vartheta\alpha])<\vartheta\beta$.
@@ 3.4. The Path Not Taken
Since we don’t actually carry out the substitutions, we no longer need to worry about variable clashes—in the term $({\color[rgb]{0,0,1}\vartheta}{\color[rgb]{0,0,1}\Omega}\#v)\Omega$, there is no danger of thinking the black $\Omega$ is bound, since it is not even inside ${\color[rgb]{0,0,1}\vartheta}$.
@@
If we followed through on this, we would run into the following obstacle—in order to compare $\vartheta\alpha\beta_{1}\cdots\beta_{n}$ with $\vartheta\alpha^{\prime}\beta^{\prime}_{1}\cdots\beta^{\prime}_{n^{\prime}}$, we would like to compare their uncollapsed versions. But the uncollapsed version of $\vartheta\alpha\beta_{1}\cdots\beta_{n}$ cannot be $\alpha[v_{i}\mapsto\beta_{i}]$—this loses the distinction between $\Omega$’s at different levels. Rather, we want $\alpha$ to be something like $\lambda v_{1}\cdots v_{n}.\alpha$. But this means, for instance, that $(\lambda v.v)\Omega$ is fundamentally different that $\Omega$ (because the former collapses to something like $(\vartheta v)\Omega$ while the latter collapses to $\vartheta\Omega$). We could accept such issues (indeed, this is roughly the same as the way we created equivalence classes to deal with proving well-foundedness), of course, but it seems that the solution would end up replicating the counting and shifting we had to deal with above instead of rescuing us from it.
[…]

### v2

@@ 3.2. Well-Foundedness
Since $\Omega^{(0)}>\Omega^{(-1)}>\cdots$, the ordinal terms are certainly not well-founded on their face. This is because our notation looks at ordinal terms “from above”, fixing $\Omega^{(0)}$ to be the highest cardinality. We should instead compare “from below”, demanding the ordinals share the same ground.
@@
Given an ordinal $\alpha$, we write $\alpha^{*}=\left\{\begin{array}[]{ll}\alpha&\text{if }\overline{\mathbb{FC}}(\alpha)=-\infty\\ \alpha^{\geq 0}_{-\overline{\mathbb{FC}}(\alpha)}&\text{otherwise}\end{array}\right.$. We may think of this as a normalized version of $\alpha$, in the sense that we have arranged pushed the cardinalities up as high as we are allowed to; in particular, $\overline{\mathbb{FC}}(\alpha^{*})\in\{-\infty,0\}$.
We could think of ordinal terms with $\overline{\mathbb{FC}}(\alpha)=0$ and $\mathbb{G}(\alpha)\geq n$ as being the terms of cardinality $\Omega_{n+1}$ by interpreting $\Omega^{(i)}$ as $\Omega_{n+i+1}$.
We will define two collections of sets $\mathrm{Acc}_{n}\subseteq M_{n}$ for $n\in\{-\infty\}\cup\mathbb{N}$.
@@
We let $M_{-\infty}$ be all ordinal terms of formal cardinality $\infty$- ∞. Given $M_{n}$, we let $\mathrm{Acc}_{n}$ be the well-founded part of $M_{n}$.
For $n\in\mathbb{N}$, given $\mathrm{Acc}_{i}$ for $i<n$, we set $M_{n}$ to be the set of ordinal terms $\alpha$ with $\overline{\mathbb{FC}}(\alpha)=0$, $\mathbb{G}(\alpha)\geq-n$, and $\{\beta^{*}\mid\beta\in K^{<0}\alpha\}\subseteq\bigcup_{i<n}\mathrm{Acc}_{i}$.
Note that when $\beta\in K^{<0}\alpha$, we always have $\overline{\mathbb{FC}}(\beta)<\overline{\mathbb{FC}}(\alpha)$ and therefore $\mathbb{G}(\beta^{*})>\mathbb{G}(\alpha)$.
[…]
Lemma 3.8.
Whenever $m\leq n$, $\alpha\in\mathrm{Acc}_{m}$, and $\beta\in\mathrm{Acc}_{n}$, we have $\alpha^{\geq 0}_{-(n-m)}\#\beta\in\mathrm{Acc}_{n}$.
Proof.
We proceed first by induction on $m$, then on $\alpha$, and then by induction on $\beta$. From the definitions, it is easy to see that $\alpha^{\geq 0}_{-(n-m)}\#\beta\in M_{n}$.
It suffices to show that, for any $\gamma<\alpha^{\geq 0}_{-(n-m)}\#\beta$ in $M_{n}$, $\gamma\in\mathrm{Acc}_{n}$.
So let $\gamma<\alpha^{\geq 0}_{-(n-m)}\#\beta$ be given. We consider cases for $\gamma$. If $\gamma\in\mathbb{H}$ then either $\gamma\leq\alpha^{\geq 0}_{-(n-m)}$ or $\gamma\leq\beta$. In the latter case, $\gamma\in\mathrm{Acc}_{n}$ since $\beta$ is. If $\gamma\leq\alpha^{\geq 0}_{-(n-m)}$ then, since $\overline{\mathbb{FC}}(\gamma)\in\{-\infty,0\}$, we must have $m=n$, so $\alpha^{\geq 0}_{-(n-m)}=\alpha$, so $\gamma\in\mathrm{Acc}_{n}$.
In the remaining case, $\gamma=\#\{\gamma_{i}\}$. If $\gamma\leq\beta$ then we immediately have $\gamma\in\mathrm{Acc}_{n}$, so suppose $\beta<\gamma<\alpha^{\geq 0}_{-(n-m)}\#\beta$. Let $\gamma^{\prime}=\#\{\gamma_{i}\mid\alpha^{\geq 0}_{-(n-m)}\leq\gamma_{i}\}$—that is, the sum where we remove all components $<\alpha^{\geq 0}_{-(n-m)}$.
We have $\gamma^{\prime}\leq\gamma<\alpha^{\geq 0}_{-(n-m)}\#\beta$. Suppose $\overline{\mathbb{FC}}(\gamma^{\prime})=\overline{\mathbb{FC}}(\gamma)$. Then, since $\gamma^{\prime}\leq\beta$, we have $\gamma^{\prime}\in\mathrm{Acc}_{n}$. We may then repeatedly apply the main inductive hypothesis to each $\gamma_{i}<\alpha^{\geq 0}_{-(n-m)}$ to show that $\gamma\in\mathrm{Acc}_{n}$ as well.
If $\overline{\mathbb{FC}}(\gamma^{\prime})<\overline{\mathbb{FC}}(\gamma)$ then we must have $m=n$ and $\gamma^{\prime}=0$ and the result follows from $\alpha^{\geq 0}_{-(n-m)}=\alpha\in\mathrm{Acc}_{n}$ and the main inductive hypothesis. ∎
Lemma 3.9.
If $\alpha\in\mathrm{Acc}_{n}$ then $\omega^{\alpha}\in\mathrm{Acc}_{n}$.
Proof.
By induction on $n$ and then on $\alpha$, it suffices to show that for all $\beta<\omega^{\alpha}$ with $\beta\in M_{n}$, $\beta\in\mathrm{Acc}_{n}$.
We proceed by structural induction. If $\beta=\#\{\beta_{i}\}$ then each $\beta_{i}<\omega^{\alpha}$ and therefore each $\beta_{i}\in\mathrm{Acc}_{n}$ (by the side inductive hypothesis if $\overline{FC}(\beta_{i})=n$ and by the definition of $M_{n}$ and the main inductive hypothesis if $\overline{FC}(\beta_{i})<n$). Since $\bigcup_{i}\mathrm{Acc}_{i}$ is closed under sums by th eprevious lemma, $\beta\in\mathrm{Acc}_{n}$.
If $\beta=\omega^{\beta^{\prime}}$ then $\beta^{\prime}<\alpha$ and $\beta\in\mathrm{Acc}_{n}$ by the main inductive hypothesis.
If $\beta=\vartheta\beta^{\prime}$ then $\vartheta\beta^{\prime}\leq\alpha$ and therefore $\vartheta\beta^{\prime}\in\mathrm{Acc}_{n}$. ∎
Lemma 3.10.
If $\alpha\in\mathrm{Acc}_{n}$ then $\vartheta(\alpha^{\leq 0}_{-1})\in\mathrm{Acc}_{n}$.
Proof.
This includes the case where $n=-\infty$ (in which case $\alpha^{\leq 0}_{-1}=\alpha$)—this is the case where we do not actually bind any copies of $\Omega$.
We proceed by main induction on $\alpha$. It suffices to show that, for all $\gamma<\vartheta\alpha^{\geq 0}_{-1}$ with $\gamma\in M_{n}$, $\gamma\in\mathrm{Acc}_{n}$. We show this by a structural induction on $\gamma$.
If $\gamma$ is $\Omega^{(0)}$ then this is immediate, since in this case $\Omega^{(0)}$ is the smallest element of $M_{n}$. If $\gamma=\#\{\gamma_{i}\}$ then each $\gamma_{i}$ in $M_{n}$ is less than $\vartheta\alpha^{\geq 0}_{-1}$ and thefore in $\mathrm{Acc}_{n}$. If $\gamma_{i}$ is in $M_{m}$ for some $m<n$ then, since $\gamma\in M_{n}$, we always have $\gamma_{i}\in\mathrm{Acc}_{m}$. Then by the lemma above, $\gamma\in\mathrm{Acc}_{n}$.
If $\gamma=\omega^{\gamma^{\prime}}$ then $\gamma^{\prime}<\vartheta\alpha^{\geq 0}_{-1}$, so $\gamma^{\prime}\in\mathrm{Acc}_{n}$, so $\gamma\in\mathrm{Acc}_{n}$.
So suppose $\gamma=\vartheta\beta$. If $\beta<\alpha^{\geq 0}_{-1}$ then $\overline{\mathbb{FC}}(\beta)<0$, so $\beta$ is precisely of the form $\delta^{\geq 0}_{-1}$ for some $\delta$ with $\overline{\mathbb{FC}}^{\geq 0}(\delta)\in\{-\infty,0\}$. We wish to show that $\delta\in M_{n}$, so consider some $\delta^{\prime}\in K^{<0}\delta$. Then we have $(\delta^{\prime})^{\geq 0}_{-1}\in K^{<0}\beta$, so $(\delta^{\prime})^{\geq 0}_{-1}<\vartheta\alpha^{\geq 0}_{-1}$. If $\overline{\mathbb{FC}}^{\geq 0}(\delta^{\prime})=\overline{\mathbb{FC}}(\gamma)$ then the inductive hypothesis applies and $((\delta^{\prime})^{\geq 0}_{-1})^{*}\in\mathrm{Acc}_{n}$. Otherwise $\overline{\mathbb{FC}}^{\geq 0}(\delta^{\prime})<\overline{\mathbb{FC}}(\gamma)$ and, so there is $\delta^{\prime\prime}\in K^{<0}\gamma$ with $(\delta^{\prime\prime})^{*}=(\delta^{\prime})^{*}$, and so since $\gamma\in M_{n}$, $(\delta^{\prime})^{*}\in\mathrm{Acc}_{m}$ for the correct $m$. It follows that $\delta\in M_{n}$, so by the inductive hypothesis, $\gamma=\vartheta\delta^{\geq 0}_{-1}\in\mathrm{Acc}_{n}$.
Otherwise, suppose $\alpha^{\geq 0}_{-1}<\beta$. Then we must have some $\delta\in K^{<0}\alpha^{\geq 0}_{-1}$ so that $\vartheta\beta\leq\delta$. We must have $\overline{\mathbb{FC}}(\delta)<0$, and since $\gamma\in M_{n}$, it follows that $\overline{\mathbb{FC}}(\delta)=\overline{\mathbb{FC}}(\gamma)=-\infty$, so $n=-\infty$. It follows that $\delta\leq\alpha^{\geq 0}_{-1}=\alpha$, and therefore $\delta\in\mathrm{Acc}_{n}$. ∎
Lemma 3.11.
If $\alpha\in\mathrm{Acc}_{n}$ and $n>-\infty$ then $(\vartheta\alpha)^{*}\in\bigcup_{m<n}\mathrm{Acc}_{m}$.
Proof.
We focus on $\alpha$ with either $\{0,-1\}\subseteq\mathbb{FC}(\alpha)$ or $\mathbb{FC}(\alpha)=\{0\}$. (If $0\not\in\mathbb{FC}(\alpha)$, we are covered by the previous lemma. If $-1\not\in\mathbb{FC}(\alpha)$ but $\overline{\mathbb{FC}}(\vartheta\alpha)>-\infty$ then there is an $\alpha^{\prime}$ with $\{0,-1\}\subseteq\mathbb{FC}(\alpha^{\prime})$ and $(\vartheta\alpha^{\prime})^{*}=(\vartheta\alpha)^{*}$.)
We proceed by induction on $\alpha\in\mathrm{Acc}_{n}$ with this property. Since $\mathbb{G}(\vartheta\alpha)<\mathbb{G}(\alpha)$, we certainly have $(\vartheta\alpha)^{*}\in\bigcup_{m<n}M_{m}$. Let $m=\mathbb{G}((\vartheta\alpha)^{*})$.
If $\gamma$ is $\Omega^{(0)}$, this is immediate since $\Omega^{(0)}$ is the smallest element of formal cardinality 00. If $\gamma=\#\{\gamma_{i}\}$ then each $\gamma_{i}$ is either in $\mathrm{Acc}_{k}$ for some $k<m$, or $\gamma_{i}\in M_{m}$, $\gamma_{i}<(\vartheta\alpha)^{*}$, and therefore $\gamma_{i}\in\mathrm{Acc}_{m}$ by the side inductive hypothesis. Then, by closure under addition, $\gamma\in\mathrm{Acc}_{m}$.
Similarly, if $\gamma=\omega^{\gamma^{\prime}}$ then $\gamma^{\prime}<(\vartheta\alpha)^{*}$, so $\gamma^{\prime}\in\mathrm{Acc}_{m}$, and therefore $\omega^{\gamma^{\prime}}\in\mathrm{Acc}_{m}$.
For the main case, suppose $\gamma=\vartheta\gamma^{\prime}$. If $\overline{\mathbb{FC}}(\gamma^{\prime})<0$ then, by the previous lemma, $\gamma\in\mathrm{Acc}_{m}$. Otherwise, since $\vartheta\gamma^{\prime}\in M_{m}$, we must have either $\overline{\mathbb{FC}}(\vartheta\gamma^{\prime})=-\infty$ or $-1\in\mathbb{FC}(\gamma^{\prime})$. If $\delta\in K^{<0}\gamma^{\prime}$ then either $\delta\in K^{<0}\gamma$ (so $\delta^{*}\in\mathrm{Acc}_{\mathbb{G}(\delta^{*})}$ since $\gamma\in M_{m}$) or $\delta\in K^{<0}\gamma^{\prime}\setminus K^{<0}\gamma$, and therefore $\overline{\mathbb{FC}}(\delta)=-1$ and then we have $\delta<\vartheta\alpha$ and therefore, by the side inductive hypothesis, $\delta\in\mathrm{Acc}_{m}$. It follows that $\gamma^{\prime}\in M_{n}$, and therefore $\vartheta\gamma^{\prime}\in\mathrm{Acc}_{m}$ by the main inductive hypothesis.
Otherwise, there is some $\delta\in K^{<0}\alpha$ with $\gamma\leq\delta$, and since $\delta\in\mathrm{Acc}_{m}$, also $\gamma\in\mathrm{Acc}_{m}$. ∎
@@
For all $\alpha$ and all $n\geq\mathbb{G}(\alpha^{*})$, $\alpha^{*}\in\mathrm{Acc}_{n}$.
@@
By induction on terms. If $\alpha=\#\{\alpha_{i}\}$, $\alpha=\omega^{\alpha^{\prime}}$, or $\alpha=\vartheta\alpha^{\prime}$ then the claim follows from the inductive hypothesis and a lemma above. If $\alpha$ is $\Omega^{(0)}$, the claim holds because $\Omega^{(0)}$ is the smallest element of $M_{n}$. ∎
@@ 3.3. Variables
Finally, we should extend the system with a variable so that we can check the Key Lemma. We extend the notation system by variables $v^{(J)}$, where we interpret $v^{(J)}$ as a strongly critical ordinal term smaller than $\Omega^{(J)}$ and otherwise incomparable to strongly critical terms.
@@
  • $(\vartheta\alpha)[v\mapsto^{J}\beta]=\vartheta(\alpha[v\mapsto^{J-1}\beta])$,
  • $w^{(J^{\prime})}[v\mapsto^{J}\beta]=\left\{\begin{array}[]{ll}\beta^{\leq 0}_{+J}&\text{if }v=w\\ w^{(J^{\prime})}&\text{if }v\neq w\end{array}\right.$.
We should only consider substituting into variables which appear precisely at the top level—that is, we want to substitute into things like $v^{(0)}\#\vartheta v^{(1)}$ but not $v^{(1)}\#\vartheta v^{(3)}$.
@@
  • $\alpha=\vartheta\alpha^{\prime}$ and $v$ is $J-1$-substitutable in $\alpha^{\prime}$,
@@
Definition 3.15.
When $\overline{\mathbb{FC}}(\gamma)<0$, we define $D_{\gamma}(\beta)=D_{0,\gamma}(\beta)=\vartheta(\omega^{\Omega\#\beta}\#\gamma)$. We define $D_{m+1,\gamma}(\beta)=D_{0,0}(D_{m,\gamma}(\beta))$.
When $\overline{\mathbb{FC}}(\gamma)<0$, we define $\alpha\ll_{\gamma}\beta$ to hold if $\alpha<\beta$ and every $\eta\in K^{<0}\alpha$ is bounded by $D_{m,\gamma}(\beta)$ where $m$ is least so that $\overline{\mathbb{FC}}(D_{m,\gamma}(\beta))\leq\overline{\mathbb{FC}}(\eta)$.
@@
    If $\overline{\mathbb{FC}}(\delta)<0$ and $\alpha\ll_{\delta}\beta$ are ordinal terms so that every free variable is $J$-substitutable for some $J<0$ then $D_{\delta}(\alpha)\ll_{0}D_{\delta}(\beta)$.
@@
    If $\overline{\mathbb{FC}}(\delta)<0$ and $\alpha\ll_{\delta}\beta$, $\gamma\ll_{\delta}\beta$, and $v$ is 00-substitutable in $\gamma$ then $D_{D_{\delta}(\alpha)^{\leq 0}_{-1}}(\gamma[v\mapsto^{0}D_{\delta}(\alpha)^{\leq 0}_{-1}])\ll_{0}D_{\delta}(\beta)$.
Proof.
  1. (1)
    By induction on $\alpha,\beta$, following the comparison between $\alpha$ and $\beta$. Any step involving $v$ is preserved when we replace $v$ with $\gamma$.
  2. (2)
    Almost immediate from the definition.
  3. (3)
    Since $\gamma<\beta$, we also have $\gamma[v\mapsto^{0}D_{\delta}(\alpha)^{\leq 0}_{-1}]<\beta$ by (1), so $\omega^{\Omega\#\gamma[v\mapsto^{0}D_{\delta}(\alpha)^{\leq 0}_{-1}]}\#D_{\delta}(\alpha)^{\leq 0}_{-1}<\omega^{\Omega\#\beta}\#\delta$. We have $K^{<0}(\omega^{\Omega\#\gamma[v\mapsto^{0}D_{\delta}(\alpha)^{\leq 0}_{-1}]}\#D_{\delta}(\alpha)^{\leq 0}_{-1})\subseteq K^{<0}\gamma\cup K^{<0}\delta\cup\{D_{\delta}(\alpha)\}$, and so is bounded by $D_{m,\delta}(\beta)$ for the appropriate $m$.
@@ 3.4. The Path Not Taken
Since we don’t actually carry out the substitutions, we no longer need to worry about variable clashes—in the term $({\color[rgb]{0,0,1}\vartheta}{\color[rgb]{0,0,1}\Omega}\#v)(\Omega)$, there is no danger of thinking the black $\Omega$ is bound, since it is not even inside ${\color[rgb]{0,0,1}\vartheta}$.
@@
If we followed through on this, we would run into the following obstacle—in order to compare $\vartheta\alpha\beta_{1}\cdots\beta_{n}$ with $\vartheta\alpha^{\prime}\beta^{\prime}_{1}\cdots\beta^{\prime}_{n^{\prime}}$, we would like to compare their uncollapsed versions. But the uncollapsed version of $\vartheta\alpha\beta_{1}\cdots\beta_{n}$ cannot be $\alpha[v_{i}\mapsto\beta_{i}]$—this loses the distinction between $\Omega$’s at different levels. Rather, we want $\alpha$ to be something like $\lambda v_{1}\cdots v_{n}.\alpha$. But this means, for instance, that $(\lambda v.v)\Omega$ is fundamentally different that $\Omega$ (because the former collapses to something like $(\vartheta v)\Omega$ while the latter collapses to $\vartheta\Omega$). We could accept such issues, but it seems that the solution would end up replicating the counting and shifting we had to deal with above instead of rescuing us from it.
4. Functions as Cardinals